OpenStudy (anonymous):

what is 4/9 divided by 1/3

OpenStudy (anonymous):

\[\LARGE \cfrac{\cfrac ab}{\cfrac cd}=\frac{a\cdot d}{b\cdot c}\]

OpenStudy (kymber):

When dividing fractions, multiply the first fraction by the reciprocal of the second fraction!!!!!!!!!

OpenStudy (anonymous):

so 9/12

OpenStudy (kymber):

How did you get that? Can you show your work?

OpenStudy (anonymous):

i did it the other way